Tohban Report 2017-06-14

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 7 Jun 2017
End Date: 14 Jun 2017
Tohban: Milo Buitrago-Casas
Tohban email:
Next Tohban: TBD
List all reports


Solar Activity

Solar activity during last week was really low and currently is at low levels. There are two active regions on the solar disk, but solar activity is expected to be very low in the next couple of days.

How many GOES flares occurred?

 Flares above B, C, M, X class were     19     2     0     0

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were      7     1     0     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were      0     0     0     0

There were RHESSI flares/GOES flares 15 / 21 over the time range 06-Jun-17 13-Jun-17

Memory Management

The SSR was stable around 23-30% of filling. Currently is about ~25 percent.

Spacecraft Status

Detectors 1, 3, 6 and 8 are opertaing and collecting science data. Cold plate temperatures are ~155.6K and 154.0K. Cold tip temperatures are ~131 and 129K.

The Cryocooler is operating with a power of ~74.4 W and the acceleration is at ~14.6mg

Data Gaps

No data gaps this week.

Detector issues

No detector issues nor operations last week to report.

Other notes

Spacecraft is in good conditions and operating.

Spacecraft Management

A spin up was made on 23 May 2017 (2017/143 16:04 and 20:51 UTC)

Decimation Active/Vigorous
HLAT Decimation Rear decimation
Night time data (fronts) plus/minus 4 minutes from daylight
Night time data (rears) plus/minus 4 minutes from daylight
Require extra passes? No
Requirement for moving pointer? No
Attenuator operation logic enabled
Detector problems? No problems.
Personal tools